# Divergence Minimization Preference Optimization for Diffusion Model Alignment

### v1

@@ 3.3 Theoretical Analysis
In this section, we further provide the theoretical analysis of the DMPO method by showing its connection with the original RLHF objective. Specifically, we show that the optimization direction of DMPO aligns with that of maximizing the RLHF objective in Equation 3. Let the RLHF loss for diffusion path ${\mathbf{x}}_{0:T}$ be denoted as $\mathcal{L}_{\text{RLHF}}$:
@@
Starting from the divergence-based formulation in Equation 9, we formally have the following connection of the DMPO objective and the RL method objective.
@@
Theorem 2 states that the optimization direction for $p_{\theta}({\mathbf{x}}_{0:T})$ aligns with that of RLHF, which theoretically justifies that the objective of our DMPO is consistent with reinforcement learning–based alignment under specific conditions. The full technical derivation is provided in Appendix A. This highlights that though derived from a divergence minimization view, DMPO is consistent with the optimization direction of the original RLHF method.
Figure 1: We show the images generated by different models for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2. "Diff" represents "Diffusion" for simplicity.
@@ 4 Related Work
Diffusion model alignment seeks to steer generative outputs toward human preferences by embedding reinforcement learning objectives within the denoising process (Yang_2024_CVPR, ; imagereward, ; traindiffrl, ). Early efforts such as DDPO (diffrliclr, ) leverage task-specific, hand-crafted reward functions—e.g., promoting compressibility—to fine-tune pretrained diffusion backbones. More recently, DPOK (fanrl, ) replaces these bespoke rewards with feedback signals distilled from AI agents trained on vast human-preference corpora. Diffusion-DPO (diffusiondpo, ) adapts the core DPO framework, directly ingesting pairwise preference data to update the diffusion model. Diffusion-KTO kto extends utility-based preference optimization to diffusion models using only binary human feedback and DSPO (dspo, ). DSPO derives a novel step-wise alignment method for diffusion models from the perspective of score-based modeling. However, these approaches cannot precisely align the learned policy with the optimal policy derived from RLHF objective. In this work, we instead approach alignment by minimizing the divergence between the learned policy and the optimal policy, and adopt a reverse KL objective to explicitly align them. Efficient Exact Optimization (EXO) (exo, ) and $f$-divergence Preference Optimization ($f$-PO) (f-po, ) also explored aligning language models (LMs) using reverse KL. However, all these studies are limited to autoregressive language models, while our method focuses on diffusion models and provides fundamentally different inspiration for aligning the Markov chains.
@@ 5.1 Experimental Setting
Dataset and Baselines. We select Stable Diffusion v1.5 (SD1.5) and Stable Diffusion XL Base 1.0 (SDXL) as our base models. We train DMPO on the Pick-a-Pic V2 (pickapic, ) dataset, which consists of pairwise preferences for images generated by SDXL-beta and Dreamlike (a fine-tuned version of SD1.5). After excluding the $\sim$ 12% of pairs with ties, we end up with 851,293 preference pairs across 58,960 unique prompts. Our baselines include diffusion models fine-tuned on the Pick-a-Pic V2 dataset using various alignment methods, based on either SD1.5 or SDXL: Diffusion-DPO (diffusiondpo, ), Diffusion-KTO (kto, ), DSPO (dspo, ), MAPO (mapo, ). We also compare against the original SD1.5 and SDXL model as well as a supervised fine-tuning (SFT) SD1.5 model. We implement SFT and train DSPO by their released repo ourselves, and for Diffusion-DPO, KTO, and MAPO, we use the officially released checkpoints.
[…]
@@
Evaluation Details. We evaluate the performance of DPMO through both automated preference metrics and human user studies. For general alignment evaluation, we choose test prompts from Pick-a-Pic V2, HPS V2, and PartiPrompt, and compare DMPO against all baselines on text-to-image generation task. For image editing evalution, we evaluate editing performance on two standard benchmarks, TEd-bench (ted, ) and InstructPix2Pix (ip2p, ) Bench by using SDEdit (sde, ) pipeline and compare DMPO against all baselines. To ensure a fair comparison, we ensure consistency across all models, i.e., setting the guidance scale to 7.5, and the number of sampling steps to 50. For automated evaluation, we report three aspects: (1) Reward Score directly output by the reward models, (2) Win rates comparing DPMO and all baselines against the original base model, and (3) Pairwise win rates comparing DPMO directly with each baseline. The preference metrics include: PickScore (pickapic, ), HPS V2 (hpsv2, ), CLIP (clip, ), LAION Aesthetics Classifier  (laionclip, ), and ImageReward (imagereward, ), all of which are caption-aware models trained to predict human preference scores based on an image and its associated prompt. Note, In this paper, we present all DPMO and baseline results based on models trained with SD 1.5. We also conduct the same experiments and evaluations using SDXL; for more results on DPMO fine-tuning and performance on SDXL, please refer to the Appendix C.
In addition, we conduct user study to compare DPMO with existing baselines. Each comparison is assessed from three perspectives: Q1. General Preference: Which image do you prefer given the prompt? Q2. Visual Appeal: Which image is more visually appealing regardless of the prompt? and Q3. Prompt Alignment: Which image better matches the text description?. We sampled 100 test prompts from Pick-a-Pic V2, HPS V2, and PartiPrompt, and evaluated DMPO (based on SD1.5), SD1.5, and all baselines. Details on user annotations and the experimental setup is provided in the Appendix C.
@@ 5.2 Alignment Result
Qualitative Result. Figure 1 presents a qualitative comparison of generations from DMPO and other baseline models under the same prompt and control conditions. Compared to alternative methods, DMPO consistently demonstrates a stronger ability to capture the semantic intent of the prompt, producing outputs that are both more accurate and higher in quality. For instance, in the first row, only DMPO successfully renders the concept of a “smile”, while in the second row, it is the only model that correctly depicts "the papaya dressed as a sailor". Additional qualitative results can be found in the Appendix C.
@@
Quantitative Result. Tables 1 and 2 report the reward-model scores for DMPO across all the test sets. From these tables, we observe that, for every reward metric, the DMPO-tuned models consistently outperform their respective base models (SD1.5). Generally DMPO yields higher reward-model scores on Pick-a-Pic V2, HPS V2, and Parti-Prompt test set than almost all methods trained on pairwise preference data. In terms of Win rate comparison, it outperforms all other baseline models trained on Pick-a-Pic V2 by a win rate of at least 64.6% evaluated with PickScore. Moreover, Figure 2 shows the results of the user study shows human evaluators consistently prefer generations from the DMPO whether compared against the original SD1.5 or against all other baselines—further underscoring the effectiveness of the divergence minimization objective.
Pick-a-Pic V2 DMPO SD1.5 82.40 84.80 65.60 75.60 80.40
SFT SD1.5 74.60 82.60 61.60 72.40 77.80
Diff-DPO SD1.5 73.20 70.00 61.20 69.40 64.20
Diff-KTO SD1.5 73.60 84.60 58.60 74.00 80.00
DSPO SD1.5 74.60 84.80 61.60 73.60 76.20
DMPO SFT 67.20 50.60 52.80 53.20 50.20
DMPO Diff-DPO 73.40 77.20 56.80 64.80 68.40
DMPO Diff-KTO 69.00 50.80 55.20 55.20 50.20
DMPO DSPO 68.40 50.80 54.80 52.20 50.80
HPS V2 DMPO SD1.5 85.50 86.50 61.00 78.75 79.75
SFT SD1.5 77.00 89.00 56.00 73.75 81.50
Diff-DPO SD1.5 77.50 69.25 57.75 69.5 64.50
Diff-KTO SD1.5 75.25 89.50 55.00 75.75 79.25
DSPO SD1.5 74.50 89.00 58.00 78.25 78.75
DMPO SFT 67.25 47.50 56.00 57.75 53.50
DMPO Diff-DPO 74.00 74.25 58.00 70.00 72.00
DMPO Diff-KTO 68.00 47.50 56.50 55.50 53.50
DMPO DSPO 65.75 50.00 59.00 57.00 54.00
Parti Prompt DMPO SD1.5 76.72 82.67 62.01 71.69 75.25
SFT SD1.5 66.36 83.88 54.35 71.51 72.37
Diff-DPO SD1.5 67.16 64.86 54.84 60.66 63.11
Diff-KTO SD1.5 65.56 84.38 54.41 69.97 73.86
DSPO SD1.5 67.30 85.17 53.31 70.22 73.71
DMPO SFT 64.64 51.31 57.23 51.69 56.00
DMPO Diff-DPO 69.30 72.54 58.15 65.62 69.55
DMPO Diff-KTO 67.03 48.40 59.13 55.15 55.46
DMPO DSPO 66.12 47.14 59.25 53.49 55.88
@@ 5.3 Image Editing Results
Beyond improving alignment in image generation tasks, DMPO also significantly enhances the model’s capabilities in image editing, particularly for text-guided image editing scenarios. This improvement stems from the model’s strengthened ability to interpret and execute complex textual instructions. Figure 4 presents representative qualitative editing results on TEd-bench. In the first rows, only DMPO correctly understands and represents content ’pet entrance’. In the second row, where the input prompt is “A half eaten pizza.”, only DMPO generates an image that is both semantically faithful and highly visually appealing. More Qualitative Results and Quantitative results can be seen in Appendix C.
5.4 Ablation Study
Figure 3: Ablation study on parameters $\alpha$ and $\beta$, evaluated on the Pick-a-Pic V2 and HPS V2 test sets. (a) Effect of $\beta$: With $\alpha$ fixed at 0.01, model performance first increases and then decreases as $\beta$ grows. (b) Effect of $\alpha$: With $\beta$ fixed at 2000, performance also first increases and then decreases as $\alpha$ increases.
We conduct ablation studies to understand the sensitivity and behavior of our alignment objective: the effect of $\beta$ and the effect of $\alpha$. Figure 3a illustrates the performance as $\beta$ increases. As $\beta$ decreases, the optimization objective degenerates into a pure reward function, leading to a drop in performance. Conversely, as $\beta$ increases, the KL-divergence penalty becomes overly restrictive, greatly limiting the model’s capacity to adapt. Figure 3b illustrates the performance as $\alpha$ increases. Regarding the smoothing coefficient $\alpha$, we set $\alpha$ to be a small positive number since it represents the probability of the less preferred sample to avoid numerical instability. Therefore, we set $a\in\{0.1,0.01,0.001\}$ for ablation study. As shown in Appendix A, we observe that as $\alpha$ decreases, the second derivative of the objective $f(u)$ in Equation 12 increases for $u>0$, resulting in a looser upper bound. This may weaken alignment precision. In particular, when $\alpha$ is very small, the reward term becomes smoothed by the coefficient $\alpha$, causing the model to occasionally favor less-preferred samples in preference pairs.
@@ Appendix A Proofs and Derivations
A.1 Proof of Theorem 1
In this section, we provide a detailed proof of Theorem 1, demonstrating that the objective of DPO corresponds to the forward KL divergence. Starting from Equation 10 and Equation 9, we extand the dpo loss $\mathcal{L}_{dpo}(K=2)$ from pair preference data to K samples selection, where the sigmoid needs to be replaced with softmax to present the soft distribution. Utilizing the definition of $\hat{p}_{\theta}({\mathbf{x}}_{0:T})$ :
@@ A.1 Proof of Theorem 1
and by combining Equation 5, we can extend the dpo loss to:
@@
So that we can simplify the Equation 17:
@@
which completes the proof of Theorem 1.
A.2 Convexity of Equation 12
We analyze the convexity by firstly computing the second derivative of Equation 12:
@@ A.3 Proof of Theorem 2
so that we complete the proof of Theorem 2.
@@ B.1 Details of Evaluation
where $r$ is a pretrained vision–language reward model that outputs a scalar alignment score. We directly report the mean reward score over the entire test set for Table 1, Table 3 and Table 5:
Reward $\text{Reward Score}=\frac{1}{N}\sum_{i=1}^{N}R\bigl{(}I_{\text{1}}^{(i)},p^{(i)}\bigr{)}.$
@@
Each comparison is evaluated by 2 to 3 human annotators, yielding a total of 1500 judgments. For each comparison, the annotator is shown a prompt and two images generated by different models, and is asked to answer the three evaluation questions outlined in Section 5.1. We report the final win rates by aggregating the human preferences across all comparisons, as shown in Figure 2.

### v2

@@ 3.3 Theoretical Analysis
In this section, we further provide the theoretical analysis of the DMPO method by showing its connection with the original RLHF objective. Specifically, we show that the optimization direction of DMPO aligns with that of maximizing the RLHF objective in Equation˜3. Let the RLHF loss for diffusion path ${\mathbf{x}}_{0:T}$ be denoted as $\mathcal{L}_{\text{RLHF}}$:
@@
Starting from the divergence-based formulation in Equation˜9, we formally have the following connection of the DMPO objective and the RL method objective.
@@
Theorem˜2 states that the optimization direction for $p_{\theta}({\mathbf{x}}_{0:T})$ aligns with that of RLHF, which theoretically justifies that the objective of our DMPO is consistent with reinforcement learning–based alignment under specific conditions. The full technical derivation is provided in Appendix˜A. This highlights that though derived from a divergence minimization view, DMPO is consistent with the optimization direction of the original RLHF method.
Figure 2: Qualitative result of different alignment methods. We show the images generated by different models for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2. The top two rows present results based on SD1.5, while the bottom two rows are based on SDXL. "Diff" represents "Diffusion" for simplicity.
@@ 4 Related Work
Diffusion model alignment seeks to steer generative outputs toward human preferences by embedding reinforcement learning objectives within the denoising process (Yang et al., 2024a; Xu et al., 2023; Miao et al., 2024). Early efforts such as DDPO (Black et al., 2024) leverage task-specific, hand-crafted reward functions—e.g., promoting compressibility—to fine-tune pretrained diffusion backbones. More recently, DPOK (Fan et al., 2023) replaces these bespoke rewards with feedback signals distilled from AI agents trained on vast human‐preference corpora. Diffusion-DPO (Wallace et al., 2024) adapts the core DPO framework, directly ingesting pairwise preference data to update the diffusion model. Diffusion-KTO (Li et al., 2024) extends utility-based preference optimization to diffusion models using only binary human feedback and DSPO (Zhu et al., 2025). DSPO derives a novel step-wise alignment method for diffusion models from the perspective of score-based modeling. However, these approaches cannot precisely align the learned policy with the optimal policy derived from RLHF objective. In this work, we instead approach alignment by minimizing the divergence between the learned policy and the optimal policy, and adopt a reverse KL objective to explicitly align them. Efficient Exact Optimization (EXO) (Ji et al., 2024) and $f$-divergence Preference Optimization ($f$-PO) (Han et al., 2024) also explored aligning language models (LMs) using reverse KL. However, all these studies are limited to autoregressive language models, while our method focuses on diffusion models and provides fundamentally different inspiration for aligning the Markov chains.
@@ 5 Experiments
5.1 Illustrative Study of Alignment Objectives
To further illustrate the behavior of different alignment objectives, we design a toy 2D experiment. The pretraining data is sampled from a Gaussian distribution with mean $(3,3)$ and variance $2$. Preference data comes from a mixture of two Gaussians with means $(4.3,4.3)$ and $(1.8,1.8)$ (variance $1$, ratio 6:46{:}4). Undesirable data is drawn from a Gaussian with mean $(4.5,0.8)$ and variance $1$. We first pretrain a small MLP diffusion model on the pretraining data using the standard diffusion objective, and then fine-tune it with different alignment objectives. From each trained model, we sample $2000$ points. Figure 1 shows the results. DiffusionDPO pushes samples away from the undesirable region but distributes them rather uniformly, exhibiting mean-seeking behavior. In contrast, DMPO captures the dominant mode of the mixture distribution, demonstrating mode-seeking behavior.
@@ 5.2 Experimental Setting
Dataset and Baselines. We select Stable Diffusion v1.5 (SD1.5) and Stable Diffusion XL Base 1.0 (SDXL) as our base models. We train DMPO on the Pick-a-Pic V2 (Kirstain et al., 2023) dataset, which consists of pairwise preferences for images generated by SDXL-beta and Dreamlike (a fine-tuned version of SD1.5). After excluding the $\sim$ 12% of pairs with ties, we end up with 851,293 preference pairs across 58,960 unique prompts. Our baselines include diffusion models fine-tuned on the Pick-a-Pic V2 dataset using various alignment methods, based on either SD1.5 or SDXL: Diffusion-DPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024), DSPO (Zhu et al., 2025), MAPO (Hong et al., 2025). We also compare against the original SD1.5 and SDXL model as well as a supervised fine-tuning (SFT) SD1.5 model. We implement SFT and train DSPO by their released repo ourselves, and for Diffusion-DPO, KTO, and MAPO, we use the officially released checkpoints. In addition, we further compare against Stable Diffusion 3.5 Medium (SD3.5-M) (Esser et al., 2024) models fine-tuned with DPO and with our proposed DMPO objective to assess the effectiveness of these methods on a stronger backbone.
Training Details. We detail the implementation setup for DMPO in this section. Training is distributed across 4 NVIDIA A100 40GB GPUs, each processing a local batch of 2 pairs, with gradient accumulation over 256 steps to achieve the desired global batch size 2048. All models are trained at fixed square resolutions using a learning rate of $1\times 10^{-5}$, with a linear warmup spanning the first 500 steps. We select the best-performing model from training for evaluation; for DPMO with SD1.5 as the base model, we set the smoothing coefficient $\alpha=0.01$ and $\beta=2000$, for DPMO with SDXL as the base model, we use $\alpha=0.01$ and $\beta=6000$, and for DPMO with SD3.5-M as the base model, we use $\alpha=0.1$ and $\beta=100$. For more details on the model’s behavior under various parameter settings, please refer to Section˜C.4.
@@
Evaluation Details. We evaluate the performance of DPMO through both automated preference metrics and human user studies. For general alignment evaluation, we choose test prompts from Pick-a-Pic V2, HPS V2, and PartiPrompt, and compare DMPO against all baselines on text-to-image generation task. For image editing evalution, we evaluate editing performance on two standard benchmarks, TEd-bench (Kawar et al., 2023) and InstructPix2Pix (Brooks et al., 2023) Bench by using SDEdit (Meng et al., 2022) pipeline and compare DMPO against all baselines. To ensure a fair comparison, we ensure consistency across all models, i.e., setting the guidance scale to 7.5, and the number of sampling steps to 50. For automated evaluation, we report three aspects: (1) Reward Score directly output by the reward models, (2) Win rates comparing DPMO and all baselines against the original base model, and (3) Pairwise win rates comparing DPMO directly with each baseline. The preference metrics include: PickScore (Kirstain et al., 2023), HPS V2 (Wu et al., 2023), CLIP (Radford et al., 2021), LAION Aesthetics Classifier  (Schuhmann, 2022), and ImageReward (Xu et al., 2023), all of which are caption-aware models trained to predict human preference scores based on an image and its associated prompt. In addition, we conduct user study to compare DPMO with existing baselines. Each comparison is assessed from three perspectives: Q1. General Preference: Which image do you prefer given the prompt? Q2. Visual Appeal: Which image is more visually appealing regardless of the prompt? and Q3. Prompt Alignment: Which image better matches the text description?. We sampled 100 test prompts from Pick-a-Pic V2, HPS V2, and PartiPrompt, and evaluated DMPO (based on SD1.5), SD1.5, and all baselines. More details on the experimental setup is provided in the Appendix˜C.
@@ 5.3 Alignment Result
Qualitative Result. Figure˜2 presents a qualitative comparison of generations from DMPO and other baseline models under the same prompt and control conditions. Compared to alternative methods, DMPO consistently demonstrates a stronger ability to capture the semantic intent of the prompt, producing outputs that are both more accurate and higher in quality. For instance, in the second row, it is the only model that correctly depicts "the papaya dressed as a sailor". Additional qualitative results can be found in the Appendix˜C.
@@
Pick-a-Pic V2 SDXL 0.2203 0.2661 0.3609 5.9892 0.5111
DMPO 0.2264 0.2730 0.3741 5.9422 0.8563
SFT 0.2224 0.2675 0.3624 5.9239 0.5834
Diff-DPO 0.2256 0.2709 0.3722 5.9890 0.7584
MAPO 0.2213 0.2682 0.3615 6.1196 0.6226
DSPO 0.2256 0.2684 0.3615 5.9598 0.6831
HPS V2 SDXL 0.2271 0.2730 0.3775 6.1125 0.6749
DMPO 0.2320 0.2804 0.3914 6.1883 1.0169
SFT 0.2277 0.2762 0.3784 6.0638 0.6998
Diff-DPO 0.2314 0.2777 0.3890 6.1546 0.9610
MAPO 0.2279 0.2765 0.3801 6.2134 0.7839
DSPO 0.2285 0.2754 0.3795 6.0545 0.7385
Parti Prompt SDXL 0.2249 0.2714 0.3551 5.7648 0.6010
DMPO 0.2290 0.2780 0.3769 5.8598 1.0399
SFT 0.2257 0.2701 0.3531 5.7239 0.6953
Diff-DPO 0.2286 0.2763 0.3688 5.7900 0.9515
MAPO 0.2250 0.2732 0.3561 5.9089 0.7031
DSPO 0.2256 0.2714 0.3596 5.7598 0.7600
Table 3: Reward Score comparisons on Pick-a-Pic V2, HPS V2 and Parti-Prompt datasets for all baselines versus SD3.5-M, best results are in boldface.
Pick-a-Pic V2 SD3.5-M 0.2223 0.2848 0.3579 5.8989 0.9789
SD3.5-DPO 0.2281 0.2890 0.3631 6.0040 1.2241
SD3.5-DMPO 0.2294 0.2904 0.3641 6.0876 1.2456
HPS V2 SD3.5-M 0.2289 0.2917 0.3733 6.0018 1.1211
SD3.5-DPO 0.2342 0.2948 0.3722 6.1103 1.2775
SD3.5-DMPO 0.2373 0.2984 0.3762 6.1883 1.2874
Parti Prompt SD3.5-M 0.2283 0.2915 0.3620 5.6392 1.1572
SD3.5-DPO 0.2349 0.2985 0.3781 5.8215 1.4443
SD3.5-DMPO 0.2368 0.2983 0.3799 5.8598 1.4632
@@
Quantitative Result. Tables˜1, 2 and 3 report the reward-model scores for DMPO across different base models and test sets. We find that, for every reward metric, the DMPO-tuned models consistently surpass their corresponding base models. Notably, in terms of PickScore, DMPO achieves the best results across all base models and all test sets. Moreover, Tables˜4, 5 and 6 show the win rate comparisons: while DMPO generally outperforms nearly all baselines, it is particularly strong in the SD1.5-based setting, where DMPO based on SD1.5 exceeds all baseline models by at least 64.6% on every dataset. further highlighting the effectiveness of the proposed divergence minimization objective.
@@ 5.4 Image Editing Results
Beyond improving alignment in image generation tasks, DMPO also significantly enhances the model’s capabilities in image editing, particularly for text-guided image editing scenarios. This improvement stems from the model’s strengthened ability to interpret and execute complex textual instructions. Figure˜4 presents representative qualitative editing results on TEd-bench. In the first rows, only DMPO correctly understands and represents content ’hugging’. In the second row, where the input prompt is “A teddy bear holding a cup.”, only DMPO generates an image that is both semantically faithful and highly visually appealing. More results can be seen in Appendix˜C.
@@ Appendix A Proofs and Derivations
A.1 Proof of Theorem˜1
In this section, we provide a detailed proof of Theorem˜1, demonstrating that the objective of DPO corresponds to the forward KL divergence. Starting from Equation˜10 and Equation˜9, we extand the dpo loss $\mathcal{L}_{dpo}(K=2)$ from pair preference data to K samples selection, where the sigmoid needs to be replaced with softmax to present the soft distribution. Utilizing the definition of $\hat{p}_{\theta}({\mathbf{x}}_{0:T})$ :
@@ A.1 Proof of Theorem˜1
and by combining Equation˜5, we can extend the dpo loss to:
@@
So that we can simplify the Equation˜17:
@@
which completes the proof of Theorem˜1.
A.2 Convexity of Equation˜12
We analyze the convexity by firstly computing the second derivative of Equation˜12:
@@ A.3 Proof of Theorem˜2
so that we complete the proof of Theorem˜2.
@@ B.1 Details of Evaluation
where $r$ is a pretrained vision–language reward model that outputs a scalar alignment score. We directly report the mean reward score over the entire test set for Table˜1, Table˜7 and Table˜2:
$\text{Reward Score}=\frac{1}{N}\sum_{i=1}^{N}R\bigl(I_{\text{1}}^{(i)},p^{(i)}\bigr).$
@@
Each comparison is evaluated by 2 to 3 human annotators, yielding a total of 1500 judgments. For each comparison, the annotator is shown a prompt and two images generated by different models, and is asked to answer the three evaluation questions outlined in Section˜5.2. We report the final win rates by aggregating the human preferences across all comparisons, as shown in Figure˜3.